Let f(x,y)=x^3+y^3−12xy, then, Select one:
a. f has a local maximum at (4,4)

b. f has a local minimum at (0,0)

c. f has a saddle at (0,0)

d. f has a saddle at (4,4)
1
Expert's answer
2020-05-25T20:29:56-0400

"Since\ f(x,y)=x^3+y^3-12xy, \text{we have}\\
f_x=\frac{\partial f}{\partial x}=3x^2-12y,\\
f_y=\frac{\partial f}{\partial y}=3y^2-12x,\\
f_{xx}=\frac{\partial^2 f}{\partial x^2}=6x,\\
f_{yy}=\frac{\partial^2 f}{\partial y^2}=6y,\\
f_{xy}=\frac{\partial^2 f}{\partial y\partial x}=-12,\\
and\\
f_{yx}=-12.\\
Putting \ 
f_{x}=f_{y}=0, \\
\text{we obtain the critical points}\\
x=0,y =0 \ and\ x=4,y =4\\
at \ (x,y)=(0,0), \text{we get}\\
f_{xx} f_{yy}-f_{xy}^2=0-144=-144<0.\\
\text{This implies that} \\
f \ \text{has a saddle at } (0,0).\\
\text{So, the correct answer is}\ \ c"